My Teacher Gave Me This Equation Today:
5x + 6y = 3x + 2
How Would You Solve This Equation?

Answer by Alan3354(69443)

5x + 6y = 3x + 2
------------
You can simplify it, but you can't solve 1 equation in 2 unknowns.
5x + 6y = 3x + 2
2x + 6y = 2
x + 3y = 1